9. The government has decided to force firms to reduce their emissions in 3% per year. This can be achieved using the following technologies.

Technology

A B
Installation cost 12000 10000
Operating costs:
energy 2000 2100
maintenance 3000 3200
labour 800 700
insurance 240 200

Each technology has a total life of three years. The installation cost are paid at the beginning the first year (year 0). The operating cost are paid each year (year 0, 1 and 2).

a) Assuming a discount rate of 8%, calculate which is the most cost-effective technology. Write the net present value of the most cost-effective technology.

b) What is the most cost effective project?
Answer #1

9.

A

R = 8%

For Technology A,

Yearly operating cost = 2000 + 3000 + 800 + 240 = $6040

So,

Net present value of technology A = -( 12000 + 6040 + 6040/1.08 + 6040/1.08^2)

Net present value of technology A = -$28810.92 or $28811

For Technology B,

Yearly operating cost = 2100 + 3200 + 700 + 200 = $6200

Net present value of technology B = -(10000 + 6200 + 6200/1.08 + 6200/1.08^2)

Net present value of technology B = -$27256.2 or -$27256

Since the present value of cost is lower for technology B that is $27256 in comparison to that of technology A (28811). So, technology B is most cost effective.

B.

Technology B is most cost effective project.